huawfhuhdcuifcuas:

Which of the following lines in standard form is equivalent to y=2x+1?

ramen:

Can you please post the following lines? thanks

huawfhuhdcuifcuas:

ok

huawfhuhdcuifcuas:

A. 2x+y=1 B. x+2y=1 C. −2x+y=1 D. −2x−y=−1

huawfhuhdcuifcuas:

hello

jimthompson5910:

Standard form is Ax+By = C With A, B, and C as integers So to turn y = 2x+1 into standard form, you have to get that 2x term to the left side somehow